1. Dielectric Material:_____


Dielectric Constant: 1


Capacitance: 100 µF


Potential Difference: 120 V


Charge:_____


Energy:_____



2. Dielectric Material: Paper


Dielectric Constant:_____


Capacitance:_____


Potential Difference: 11 V


Charge: 6.6 × 10^–5 C


Energy:_____



1
Expert's answer
2022-03-15T10:37:00-0400

Dielectric material: air.

Charge

"q=CV=100*10^{-6}*120=0.012\:\rm C"

Energy

"W=\frac{qV}{2}=\frac{0.012*120}{2}=0.72\:\rm J"